-6x + y = -1 (1,7)
Found 2 solutions by funmath, Edwin McCravy:
Answer by funmath(2933)

-6x + y = -1 (1,7)
I assume that you are being asked if (1,7) is a point on the line. To do that, remember that the coordinates are (x,y)=(1,7). Substitute 1 in for x and 7 in for y and see if the equations's sides are equal. If they are it is, if not it's not.
-6(1)+7=-1
-6+7=-1
1=-1 Not true.
(1,7) is not on the line -6x+y=-1.
Happy Calculating!!!

Answer by Edwin McCravy(20086)

-6x + y = -1  (1,7)

You didn't give any instructions,
but I'm guessing you were probably
asked something like this: 

Answer yes or no as to whether the 
ordered pair given on the right is a 
solution to the equation given on the 
left:

    -6x + y = -1  (1,7)

To do this we substitute the first
coordinate of (1,7), which is 1, for
x and the second coordinate or (1,7),
which is 7, for y into the equation:

-6(1) + (7) = -1

Then we simplify

     -6 + 7 = -1

          1 = -1

This is false so therefore the
answer is no, the ordered pair (1,7)
is not a solution to the equation
-6x + y = -1  

-------------------------------------

Suppose you were given this instead:

-6x + y = -1  (-1,-7) 

To do thus we substitute the first
coordinate of (-1,-7), which is -1, for
x and the second coordinate or (-1,-7),
which is -7, for y into the equation:

-6(-1) + (-7) = -1

Then we simplify

      6 - 7 = -1

         -1 = -1

This is true so therefore the
answer is yes, the ordered pair (-1,-7)
is a solution to the equation
-6x + y = -1

---------------------------------------

Suppose you were given:

-6x + y = 1  (1,7)

To do this we substitute the first
coordinate of (1,7), which is 1, for
x and the second coordinate or (1,7),
which is 7, for y into the equation:

-6(1) + (7) = 1

Then we simplify

      -6 + 7 = 1

           1 = 1

This is true so therefore the
answer is yes, the ordered pair
(1,7) is a solution to the equation
-6x+y = 1


--------------------------

Suppose you were given:

   3x - 2y = 16  (-2,5)

To do this we substitute the first
coordinate of (-2,5), which is -2, 
for x and the second coordinate or
(-2,5), which is 5, for y into the
equation:

3(-2) - 2(5) = 16

Then we simplify

     -6 - 10 = 16

         -16 = 16

This is false so therefore the
answer is no, the ordered pair
(-2,5) is not a solution to the 
equation 3x - 2y = 16.

-----------------------

Edwin